OpenStudy (anonymous):

What wavelengths appear in the system's emission spectrum? The allowed energies of a quantum system are 0 eV, 5.0 eV , and 7.0 eV .

OpenStudy (anonymous):

I know : 1.00 eV = 1.60 * 10^-19 J h = 6.63 * 10^-34 m^2kg/s c = 3.00 * 10^8 m/s E = hc / wavelength

OpenStudy (anonymous):

I got 0, 4.0*10^6, 5.6*10^6 but those aren't right?